www.eurockot.com, 11/21/2017


Daily 16 (L-38)

Yesterday evening the three Swarm satellites were rearranged in the cleanrooms. The freon loaded Swarm No.1 was moved in 101A for another functional check, while Swarm No. 2 was moved in 101B for fuelling preparation. The pictures show the movement of the satellites and their rearrangement. Swarm No. 3 is undergoing a functional check today. Swarm No. 2 will be leak checked and prepared for freon loading while Swarm No.3 will be switched on.

D16_P1040410_s D16_P1040423_s